# Phase-field modeling of radiation-induced composition redistribution: An application to additively manufactured austenitic Fe-Cr-Ni

### v1

@@ 1 Introduction
Classical sharp-interface models typically employ a Dirichlet boundary (fixed at the thermal equilibrium concentration) or a flux-based boundary condition for point defect concentrations to describe an ideal sink behavior of a high-angle GB—or more realistically, of a dislocation core. These simulations are also performed in a 1D setting and are thus restricted to ideal GBs and systems with simple spatial variations in the sink density. More sophisticated models are therefore needed to describe microstructural effects and realistic GBs, especially relevant for AM austenitic SSs. Several sharp-interface models have been developed for RIS in austenitic SSs and austenitic Ni-Cr and ferritic Fe-Cr steels. Allen and Was [23] described composition-dependent vacancy migration energies for FCC Fe-Cr-Ni, and compared the temperature and dose predictions against proton irradiation data. The effect of Zr and Hf on mitigating RIS in austenitic SS via solute-vacancy trapping was described by Hackett et al. [24]. Duh et al. [25] and Field et al. [26] developed flux boundary conditions to simulate non-ideal sink behavior of symmetric tilt GBs in 304 austenitic SS. Yang et al. [27] and Nastar et al. [28] used experimental tracer diffusivity data to incorporate composition-dependent activation energies and frequency factors. They also incorporated non-ideal thermodynamic factors and preferential solute-SIA binding to more accurately capture the contributions to RIS. To describe RIS in Ni-Cr, Barnard et al. [29] employed ab initio parameters and compared the individual contributions of vacancy and SIA mechanisms. For the same system, Ozturk et al. [30] demonstrated that the effect of production bias can lead to non-monotonic “W”-shaped RIS profiles. For ferritic Fe-Cr alloys, Wharry et al. [31] showed that Cr is depleted near GBs at high temperatures as a result of Cr-vacancy exchange transport, but is enriched at low temperatures due to dominant Cr-SIA binding (the crossover in segregation occurs anywhere between 400°C to 800 °C depending on alloy composition and point defect energetics). The above models, though implemented in 1D, provide the essential framework needed to describe RIS.
@@ 2 Modeling framework
To describe the flux coupling between atoms ($\phi=1,2,\dots,K$) and point defects ($\nu=V,I$), we follow linear irreversible thermodynamics. The partial fluxes of alloy elements are given in terms of the chemical potential gradients of all components, as [43]:
$\boldsymbol{J}^{\nu}_{\phi}$ $=-\sum^{K}_{k=1}L^{\nu}_{\phi k}\nabla\left(\mu_{k}+\text{sign}(\nu)\mu_{\nu}\right),$ (1)
where $K$ is the total number of atomic components, $\text{sign}(V)=-1$ for the vacancy ($V$) and $\text{sign}(I)=+1$ for the SIA ($I$, which can be any atom $\phi$). Due to conservation, the total fluxes are given by $\boldsymbol{J}_{\phi}=\sum_{\nu}J^{\nu}_{\phi}$ and $\boldsymbol{J}_{\nu}=\text{sign}(\nu)\sum^{K}_{k=1}\boldsymbol{J}^{\nu}_{k}$. From the Onsager reciprocal relations, the rest of the transport coefficients are given by $L_{\phi\nu}=\text{sign}(\nu)\sum^{K}_{k=1}L^{\nu}_{\phi k}$ and $L_{\nu\nu}=\sum^{K}_{k=1}\sum^{K}_{l=1}L^{\nu}_{kl}$. Similar to Piochaud et al.’s treatment for a binary alloy [35], we reformulate the total fluxes in terms of the diffusion potentials ($\mu_{k1}=\mu_{k}-\mu_{1}$) by invoking the Gibbs-Duhem relation $\sum^{K}_{k=1}c_{k}\nabla\mu_{k}+\sum_{\nu}c_{\nu}\nabla\mu_{\nu}=0$ and considering the point defect concentrations in their dilute limits. Here, the lattice site concentrations are related as $c_{1}=1-\sum^{K}_{k=2}c_{k}+\sum_{\nu}\text{sign}(\nu)c_{\nu}\approx 1-\sum^{K}_{k=2}c_{k}$, where $1$ refers to the solvent atom, with the rest being solute atoms (here, $c_{k}$ counts both the on-lattice atoms and the off-lattice SIA atoms). Thus (see Sec. S1 in the Supplementary Material):
$\boldsymbol{J}_{\phi}$ $=-\sum^{K}_{k=2}L^{1}_{\phi k}\nabla\mu_{k1}-\sum_{\nu=V,I}L_{\phi\nu}\nabla\mu_{\nu},$ (2a)
$\boldsymbol{J}_{\nu}$ $=-\sum^{K}_{k=2}L^{1}_{\nu k}\nabla\mu_{k1}-L_{\nu\nu}\nabla\mu_{\nu},$ (2b)
@@
$L^{1}_{\nu k}=\text{sign}(\nu)\left(\sum^{K}_{j=1}L^{\nu}_{kj}-c_{k}L_{\nu\nu}\right),$ (3a)
$L^{1}_{\phi k}=\sum_{\nu}L^{\nu}_{\phi k}-\,c_{k}\sum_{\nu}\text{sign}(\nu)L_{\phi\nu}.$ (3b)
@@
$\frac{\partial c_{\nu}}{\partial t}=-\nabla\cdot\boldsymbol{J}_{\nu}+P_{\nu}-R_{VI}c_{V}c_{I}-k^{2}_{\nu}D_{\nu}(c_{\nu}-c^{e}_{\nu}),$ (4b)
where $D_{\nu}=L_{\nu\nu}\theta_{\nu\nu}$ represent the point defect diffusivities. In Eq. 4b, the non-conserved mean-field rate terms are the point defect production rate $P_{\nu}$, recombination reaction rate $R_{VI}$, and sink strength $k^{2}_{\nu}$. Coupling of the above with the PF microstructure for polycrystals and dislocation cells is presented in the subsequent sections.
@@ 2.1 PF model for GB segregation
We adapt the grand potential approach originally proposed for binary alloy solidification by Plapp [44], along with the extension to multicomponent and mutiphase systems, as proposed by Aagesen et al. [45]. The isothermal grand potential functional for the polycrystalline microstructure is written as:
@@
where the grand potential densities for the bulk ($b$) and GB ($g$) are given by $\omega^{\psi}=f^{\psi}-\sum^{K}_{k=2}c^{\psi}_{k}\mu_{k1}-\sum_{\nu}c_{\nu}\mu_{\nu}$ ($\psi=b$ or $g$). $\omega^{g}-\omega^{b}$ is the excess grand potential density in the GB region, $m_{0}$ is an external barrier height for the GB, and $\kappa$ is the gradient energy coefficient for the order parameter fields. For simplicity, we only consider distinct energetics (at the GB relative to the bulk) for the alloy component, and ignore distinct energetics of the point defects. While the second term in the functional penalizes large GB widths, the last term penalizes small GB widths. We only consider gradient energy contributions for the order parameter fields, and ignore their contributions for the concentration fields [45]. We use a normalized multiwell potential given by $\bar{g}_{\text{mw}}=8g_{\text{mw}}$ (see Fig. 1), where $g_{\text{mw}}$ is the standard multiwell potential employed in existing PF models of grain growth [46, 45]:
@@
$+\sum^{K}_{k=2}\sum_{l\neq k}\frac{\theta^{\psi,\circ}_{kl}}{2}(c^{\psi}_{k}-c^{\psi,\circ}_{k})(c^{\psi}_{l}-c^{\psi,\circ}_{l})+\frac{RT}{V_{m}}\sum_{\nu}c_{\nu}\ln{\left(\frac{c_{\nu}}{c^{e}_{\nu}}\right)}.$ (7)
@@
$c_{\nu}$ $=-\frac{\delta\Omega}{\delta\mu_{\nu}}=-\frac{\partial\omega^{b}}{\partial\mu_{\nu}}=-\frac{\partial\omega^{g}}{\partial\mu_{\nu}}.$ (8b)
By inverting the expressions $\mu_{\phi 1}=\partial f/\partial c_{\phi}$ and $\mu_{\nu}=\partial f/\partial c_{\nu}$ obtained from Eq. 2.1, the phase concentrations can be expressed in terms of the diffusion potentials as:
@@
$c^{\psi}_{\nu}$ $=c_{\nu}=c^{e}_{\nu}\exp{\left(\frac{V_{m}}{RT}\mu_{\nu}\right)},$ (9b)
@@
$\frac{\partial c_{\phi}}{\partial t}=\nabla\cdot\left(\sum^{K}_{k=2}L^{1}_{\phi k}\nabla\mu_{k1}+\sum_{\nu=V,I}L_{\phi\nu}\nabla\mu_{\nu}\right),$ (10a)
$\frac{\partial c_{\nu}}{\partial t}=\nabla\cdot\left(\sum^{K}_{k=2}L^{1}_{\nu k}\nabla\mu_{k1}+L_{\nu\nu}\nabla\mu_{\nu}\right),$ (10b)
$\hskip 28.45274pt+P_{\nu}-R_{VI}c_{V}c_{I}-k^{2}_{\nu,b}D_{\nu}(c_{\nu}-c^{e}_{\nu})-k^{2}_{g}D_{\nu}(c_{\nu}-c^{e}_{\nu})g_{\text{sink}}(\boldsymbol{\eta}),$ (10c)
where $k^{2}_{\nu,b}$ is the mean-field sink strength of dislocations in the bulk and is distinct for vacancies and SIAs. $k^{2}_{g}$ is the local sink strength of the GB acting in the GB region that is identified by $g_{\text{sink}}$. Thus, $k^{2}_{g}$ must be determined in conjunction with $g_{\text{sink}}$. In A and B, we test three different functions (see Fig. 1) for $g_{\text{sink}}$ and compare the results of point defect and atomic concentrations from the PF implementation against that of the sharp-interface implementations of GB sinks. Based on these results, we propose $g_{\text{sink}}$ as:
@@
$\frac{\partial c_{\phi}}{\partial t}=$ $\nabla\cdot\left(\sum^{K}_{k=2}\sum^{K}_{j=2}L^{1}_{\phi j}\theta_{jk}\nabla c_{k}+\sum_{\nu=V,I}L_{\phi\nu}\theta_{\nu\nu}\nabla c_{\nu}\right.$
@@
$\frac{\partial c_{\nu}}{\partial t}=$ $\nabla\cdot\left(\sum^{K}_{k=2}\sum^{K}_{j=2}L^{1}_{\nu j}\theta_{jk}\nabla c_{k}+L_{\nu\nu}\theta_{\nu\nu}\nabla c_{\nu}\right)$
$+P_{\nu}-R_{VI}c_{V}c_{I}-k^{2}_{\nu,b}D_{\nu}(c_{\nu}-c^{e}_{\nu})-k^{2}_{g}D_{\nu}(c_{\nu}-c^{e}_{\nu})g_{\text{sink}}(\boldsymbol{\eta}).$ (12b)
@@ 2.2 Spatially resolved model for CW segregation
$\frac{\partial c_{\phi}}{\partial t}=$ $\nabla\cdot\left(\sum^{K}_{k=2}\sum^{K}_{j=2}L^{1}_{\phi j}\theta_{jk}\nabla c_{k}+\sum_{\nu=V,I}L_{\phi\nu}\theta_{\nu\nu}\nabla c_{\nu}\right),$ (14a)
$\frac{\partial c_{\nu}}{\partial t}=$ $\nabla\cdot\left(\sum^{K}_{k=2}\sum^{K}_{j=2}L^{1}_{\nu j}\theta_{jk}\nabla c_{k}+L_{\nu\nu}\theta_{\nu\nu}\nabla c_{\nu}\right)$
$+P_{\nu}-R_{VI}c_{V}c_{I}-k^{2}_{\nu,b}(\rho)D_{\nu}(c_{\nu}-c^{e}_{\nu}),$ (14b)
where $k^{2}_{\nu,b}(\rho)$ is the spatially varying sink strength. The distribution in dislocation density and initial concentration (AM microsegregation) are defined as:

### v2

# Sourabh B. Kadambi Daniel Schwen Jia-Hong Ke Lingfeng He Andrea Jokisaari

@@ 1 Introduction
Classical sharp-interface models typically employ a Dirichlet boundary (fixed at the thermal equilibrium concentration) to describe an ideal sink behavior of a high-angle GB—or more realistically, of a dislocation core. These simulations are also performed in a 1D setting and are thus restricted to ideal GBs and systems with simple spatial variations in the sink density. More sophisticated models are therefore needed to describe microstructural effects and realistic GBs, especially relevant for AM austenitic SSs. Several sharp-interface models have been developed for RIS in austenitic SSs and austenitic Ni-Cr and ferritic Fe-Cr steels. Allen and Was [23] described composition-dependent vacancy migration energies for FCC Fe-Cr-Ni, and compared the temperature and dose predictions against proton irradiation data. The effect of Zr and Hf on mitigating RIS in austenitic SS via solute-vacancy trapping was described by Hackett et al. [24]. Duh et al. [25] and Field et al. [26] developed flux boundary conditions to simulate non-ideal sink behavior of symmetric tilt GBs in 304 austenitic SS. Yang et al. [27] and Nastar et al. [28] used experimental tracer diffusivity data to incorporate composition-dependent activation energies and frequency factors. They also incorporated non-ideal thermodynamic factors and preferential solute-SIA binding to more accurately capture the contributions to RIS. To describe RIS in Ni-Cr, Barnard et al. [29] employed ab initio parameters and compared the individual contributions of vacancy and SIA mechanisms. For the same system, Ozturk et al. [30] demonstrated that the effect of production bias can lead to non-monotonic “W”-shaped RIS profiles. For ferritic Fe-Cr alloys, Wharry et al. [31] showed that Cr is depleted near GBs at high temperatures as a result of Cr-vacancy exchange transport, but is enriched at low temperatures due to dominant Cr-SIA binding (the crossover in segregation occurs anywhere between 400°C to 800 °C depending on alloy composition and point defect energetics). The above models, though implemented in 1D, provide the essential framework needed to describe RIS.
@@ 2 Modeling framework
To describe the flux coupling between atoms ($\phi=1,2,\dots,K$) and point defects (${{\upsilon}}=V,I$), we follow linear irreversible thermodynamics. The partial fluxes of alloy elements are given in terms of the chemical potential gradients of all components, as [43]:
$\boldsymbol{J}^{{\upsilon}}_{\phi}$ $=-\sum^{K}_{k=1}L^{{\upsilon}}_{\phi k}\nabla\left(\mu_{k}+\text{sign}({{\upsilon}})\mu_{{\upsilon}}\right),$ (1)
where $K$ is the total number of atomic components, $\text{sign}(V)=-1$ for the vacancy ($V$) and $\text{sign}(I)=+1$ for the SIA ($I$, which can be any atom $\phi$). Due to conservation, the total fluxes are given by $\boldsymbol{J}_{\phi}=\sum_{{\upsilon}}J^{{\upsilon}}_{\phi}$ and $\boldsymbol{J}_{{\upsilon}}=\text{sign}({{\upsilon}})\sum^{K}_{k=1}\boldsymbol{J}^{{\upsilon}}_{k}$. From the Onsager reciprocal relations, the rest of the transport coefficients are given by $L_{\phi{{\upsilon}}}=\text{sign}({{\upsilon}})\sum^{K}_{k=1}L^{{\upsilon}}_{\phi k}$ and $L_{{{\upsilon}}{{\upsilon}}}=\sum^{K}_{k=1}\sum^{K}_{l=1}L^{{\upsilon}}_{kl}$. Similar to Piochaud et al.’s treatment for a binary alloy [35], we reformulate the total fluxes in terms of the diffusion potentials ($\mu_{k1}=\mu_{k}-\mu_{1}$) by invoking the Gibbs-Duhem relation $\sum^{K}_{k=1}c_{k}\nabla\mu_{k}+\sum_{{\upsilon}}c_{{\upsilon}}\nabla\mu_{{\upsilon}}=0$ and considering the point defect concentrations in their dilute limits. Here, the lattice site concentrations are related as $c_{1}=1-\sum^{K}_{k=2}c_{k}+\sum_{{\upsilon}}\text{sign}({{\upsilon}})c_{{\upsilon}}\approx 1-\sum^{K}_{k=2}c_{k}$, where $1$ refers to the solvent atom, with the rest being solute atoms (here, $c_{k}$ counts both the on-lattice atoms and the off-lattice SIA atoms). Thus (see Sec. S1 in the Supplementary Material):
$\boldsymbol{J}_{\phi}$ $=-\sum^{K}_{k=2}L^{1}_{\phi k}\nabla\mu_{k1}-\sum_{{{\upsilon}}=V,I}L_{\phi{{\upsilon}}}\nabla\mu_{{\upsilon}},$ (2a)
$\boldsymbol{J}_{{{\upsilon}}}$ $=-\sum^{K}_{k=2}L^{1}_{{{\upsilon}}k}\nabla\mu_{k1}-L_{{{\upsilon}}{{\upsilon}}}\nabla\mu_{{\upsilon}},$ (2b)
@@
$L^{1}_{{{\upsilon}}k}=\text{sign}({{\upsilon}})\left(\sum^{K}_{j=1}L^{{\upsilon}}_{kj}-c_{k}L_{{{\upsilon}}{{\upsilon}}}\right),$ (3a)
$L^{1}_{\phi k}=\sum_{{{\upsilon}}}L^{{\upsilon}}_{\phi k}-\,c_{k}\sum_{{{\upsilon}}}\text{sign}({{\upsilon}})L_{\phi{{\upsilon}}}.$ (3b)
@@
$\frac{\partial c_{{\upsilon}}}{\partial t}=-\nabla\cdot\boldsymbol{J}_{{\upsilon}}+P_{{\upsilon}}-R_{VI}c_{V}c_{I}-k^{2}_{{\upsilon}}D_{{\upsilon}}(c_{{\upsilon}}-c^{e}_{{\upsilon}}),$ (4b)
where $D_{{\upsilon}}=L_{{{\upsilon}}{{\upsilon}}}\theta_{{{\upsilon}}{{\upsilon}}}$ represent the point defect diffusivities, with $\theta_{\upsilon\upsilon}$ being the thermodynamic factor. $c^{e}_{\upsilon}=\exp{\left(\frac{-E^{f}_{\upsilon}+S^{f}_{\upsilon}T}{k_{B}T}\right)}$ is the point defect concentration at thermal equilibrium, with $E^{f}_{\upsilon}$ being the formation energy and $S^{f}_{\upsilon}$ being the formation entropy. In Eq. 4b, the non-conserved mean-field rate terms are the point defect production rate $P_{{\upsilon}}$, recombination reaction rate $R_{VI}$, and sink strength $k^{2}_{{\upsilon}}$. Coupling of the above with the PF microstructure for polycrystals and dislocation cells is presented in the subsequent sections.
@@ 2.1 PF model for GB segregation
We adapt the grand potential approach originally proposed for binary alloy solidification by Plapp [44], along with the extension to multicomponent and multiphase systems, as proposed by Aagesen et al. [45]. The isothermal grand potential functional for the polycrystalline microstructure is written as:
@@
where the grand potential densities for the bulk ($b$) and GB ($g$) are given by $\omega^{\psi}=f^{\psi}-\sum^{K}_{k=2}c^{\psi}_{k}\mu_{k1}-\sum_{{\upsilon}}c_{{\upsilon}}\mu_{{\upsilon}}$ ($\psi=b$ or $g$). $\omega^{g}-\omega^{b}$ is the excess grand potential density in the GB region, $\bar{g}_{\text{mw}}$ is a normalized multiwell potential, $m_{0}$ is an external barrier height for the GB, and $\kappa$ is the gradient energy coefficient for the order parameter fields. For simplicity, we only consider distinct energetics (at the GB relative to the bulk) for the alloy component, and ignore distinct energetics of the point defects. While the second term in the functional penalizes large GB widths, the last term penalizes small GB widths. We only consider gradient energy contributions for the order parameter fields, and ignore their contributions for the concentration fields [45]. $\bar{g}_{\text{mw}}$ is given by $8g_{\text{mw}}$ (see Fig. 1), where $g_{\text{mw}}$ is the standard multiwell potential employed in existing PF models of grain growth [46, 45]:
@@
$+\sum^{K}_{k=2}\sum_{l\neq k}\frac{\theta^{\psi,\circ}_{kl}}{2}(c^{\psi}_{k}-c^{\psi,\circ}_{k})(c^{\psi}_{l}-c^{\psi,\circ}_{l})+\frac{RT}{V_{m}}\sum_{{\upsilon}}c_{{\upsilon}}\ln{\left(\frac{c_{{\upsilon}}}{c^{e}_{{\upsilon}}}\right)}.$ (7)
@@
$c_{{{\upsilon}}}$ $=-\frac{\delta\Omega}{\delta\mu_{{{\upsilon}}}}=-\frac{\partial\omega^{b}}{\partial\mu_{{\upsilon}}}=-\frac{\partial\omega^{g}}{\partial\mu_{{\upsilon}}}.$ (8b)
By inverting the expressions $\mu_{\phi 1}=\partial f/\partial c_{\phi}$ and $\mu_{{{\upsilon}}}=\partial f/\partial c_{{\upsilon}}$ obtained from Eq. 2.1, the phase concentrations can be expressed in terms of the diffusion potentials as:
@@
$c^{\psi}_{{{\upsilon}}}$ $=c_{{\upsilon}}=c^{e}_{{{\upsilon}}}\exp{\left(\frac{V_{m}}{RT}\mu_{{\upsilon}}\right)},$ (9b)
@@
$\frac{\partial c_{\phi}}{\partial t}=\nabla\cdot\left(\sum^{K}_{k=2}L^{1}_{\phi k}\nabla\mu_{k1}+\sum_{{{\upsilon}}=V,I}L_{\phi{{\upsilon}}}\nabla\mu_{{\upsilon}}\right),$ (10a)
$\frac{\partial c_{{\upsilon}}}{\partial t}=\nabla\cdot\left(\sum^{K}_{k=2}L^{1}_{{{\upsilon}}k}\nabla\mu_{k1}+L_{{{\upsilon}}{{\upsilon}}}\nabla\mu_{{\upsilon}}\right),$ (10b)
$\hskip 28.45274pt+P_{{\upsilon}}-R_{VI}c_{V}c_{I}-k^{2}_{{{\upsilon}},b}D_{{\upsilon}}(c_{{\upsilon}}-c^{e}_{{\upsilon}})-k^{2}_{g}D_{{\upsilon}}(c_{{\upsilon}}-c^{e}_{{\upsilon}})g_{\text{sink}}(\boldsymbol{\eta}),$ (10c)
where $k^{2}_{{{\upsilon}},b}$ is the mean-field sink strength of dislocations in the bulk and is distinct for vacancies and SIAs. $k^{2}_{g}$ is the local sink strength of the GB acting in the GB region that is identified by $g_{\text{sink}}$. Thus, $k^{2}_{g}$ must be determined in conjunction with $g_{\text{sink}}$. In A and B, we test three different functions (see Fig. 1) for $g_{\text{sink}}$ and compare the results of point defect and atomic concentrations from the PF implementation against that of the sharp-interface implementations of GB sinks. Based on these results, we propose $g_{\text{sink}}$ as:
@@
$\frac{\partial c_{\phi}}{\partial t}=$ $\nabla\cdot\left(\sum^{K}_{k=2}\sum^{K}_{j=2}L^{1}_{\phi j}\theta_{jk}\nabla c_{k}+\sum_{{{\upsilon}}=V,I}L_{\phi{{\upsilon}}}\theta_{{{\upsilon}}{{\upsilon}}}\nabla c_{{\upsilon}}\right.$
@@
$\frac{\partial c_{{\upsilon}}}{\partial t}=$ $\nabla\cdot\left(\sum^{K}_{k=2}\sum^{K}_{j=2}L^{1}_{{{\upsilon}}j}\theta_{jk}\nabla c_{k}+L_{{{\upsilon}}{{\upsilon}}}\theta_{{{\upsilon}}{{\upsilon}}}\nabla c_{{\upsilon}}\right)$
$+P_{{\upsilon}}-R_{VI}c_{V}c_{I}-k^{2}_{{{\upsilon}},b}D_{{\upsilon}}(c_{{\upsilon}}-c^{e}_{{\upsilon}})-k^{2}_{g}D_{{\upsilon}}(c_{{\upsilon}}-c^{e}_{{\upsilon}})g_{\text{sink}}(\boldsymbol{\eta}).$ (12b)
@@ 2.2 Spatially resolved model for CW segregation
$\frac{\partial c_{\phi}}{\partial t}=$ $\nabla\cdot\left(\sum^{K}_{k=2}\sum^{K}_{j=2}L^{1}_{\phi j}\theta_{jk}\nabla c_{k}+\sum_{{{\upsilon}}=V,I}L_{\phi{{\upsilon}}}\theta_{{{\upsilon}}{{\upsilon}}}\nabla c_{{\upsilon}}\right),$ (14a)
$\frac{\partial c_{{\upsilon}}}{\partial t}=$ $\nabla\cdot\left(\sum^{K}_{k=2}\sum^{K}_{j=2}L^{1}_{{{\upsilon}}j}\theta_{jk}\nabla c_{k}+L_{{{\upsilon}}{{\upsilon}}}\theta_{{{\upsilon}}{{\upsilon}}}\nabla c_{{\upsilon}}\right)$
$+P_{{\upsilon}}-R_{VI}c_{V}c_{I}-k^{2}_{{{\upsilon}},b}(\rho)D_{{\upsilon}}(c_{{\upsilon}}-c^{e}_{{\upsilon}}),$ (14b)
where $k^{2}_{{{\upsilon}},b}(\rho)$ is the spatially varying sink strength. The distribution in dislocation density and initial concentration (AM microsegregation) are defined as: